solve for x and y


x^2+2xsin(xy)+1 = 0



Grade:

1 Answers

Kaushik Kumar
11 years ago

x^2+2xsin(xy)+1=0

sin(xy)=-(x+1/x)/2 (a number and its reciprocal always greater that or equal to 2)

therefore sin(xy)=-1

xy=-pie/2 or =3/2*pie.....

infinete solutions...

(-1, pie/2) , (1,-pie/2)......